Circular No. 7219

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999da IN NGC 6411 A. V. Filippenko, University of California at Berkeley (UCB), reports

that a CCD spectrum (range 330–1000 nm) obtained on July 9 UT with the 3-m Shane reflector at Lick Observatory reveals SN 1999da (cf. IAUC 7215) to be of type Ia, but similar to the peculiar, subluminous type-Ia SNe 1986G or 1991bg (e.g., Filippenko et al. 1992, A.J. 104, 1543). SN 1999da also appears to be subluminous, and its host is an elliptical galaxy, consistent with previous indications that such type-Ia events tend to occur in early-type galaxies.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics (CfA), report that a spectrum of SN 1999da taken by M. Calkins on July 9.4 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph) also reveals it to be type-Ia supernova that is likely subluminous. The spectrum exhibits absorption from Ti II as well as a strong Si II 580-nm feature typically seen in subluminous events such as SN 1998de (cf. IAUC 6980). Cross-correlation of the host-galaxy spectrum with an absorption-line template yields a recession velocity of 3830 ± 40 km/s for NGC 6411.

GX 17+2 P. J. Callanan, University College, Cork; A. V. Filippenko, UCB; and

M. R. Garcia, CfA, write: “We have discovered a highly variable K-band counterpart to the persistently bright x-ray burster GX 17+2. Observations obtained with with the Near Infrared Camera on Keck I during June 26.47 UT clearly resolve star ‘TR’ (Tarenghi and Reina 1972, Nat. Phys. Sci. 240, 53; Naylor et al. 1991, MNRAS 252, 203) into two components, displaced in δ by 0′′.9. The southerly and northerly components have K-band magnitudes of 14.5 and 14.9, respectively. The position of the northerly component agrees well with the astrometry of Deutsch et

al. (1999, http://xxx.lanl.gov/abs/astro-ph/?9905174), and in par-

ticular with their ‘star A’ identified with HST/NICMOS with an H-band magnitude of ∼ 19.8. For an H-band reddening of ∼ 2 mag (ibid.), and an H −K color range of 0 to +0.3 (early O–late M), this implies a K mag of ∼ 18.5–18.8 during their observations, and hence an amplitude of variability of ∼ 3.5–4 mag at K. If our counterpart is unrelated to ‘star A’ of Deutsch et al., then the amplitude of variability is even higher. In any case, this large degree of variability is surprising, considering the persistently bright nature of the x-ray emission of GX 17+2.”

1999 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green